OpenStudy (lgbasallote):

Consider the following reaction: 2A + B -> 3C + D 3.0 mol A and 2.0 mol B react to form 4.0 mol C. What is the percent yield of this reaction? A) 50% B) 67% C) 75% D) 89% E) 100% are there any other formulas for % yield other than actual yield/theoretical yield?

OpenStudy (lgbasallote):

@UnkleRhaukus how do i find the limiting reagent?

OpenStudy (lgbasallote):

lol nevermind...found it

OpenStudy (lgbasallote):

@mathslover it isnt

mathslover (mathslover):

I don't know whether the link is right or not

OpenStudy (lgbasallote):

you're not yahoo. stop insisting

OpenStudy (lgbasallote):

btw...you NEVER answered my question dont you take into consideration other factors other than just moles in deciding limiting reagent?

OpenStudy (unklerhaukus):

the limiting reagent is A

mathslover (mathslover):

r u saying that to me @lgbasallote ?

OpenStudy (lgbasallote):

to yahoo

OpenStudy (lgbasallote):

that was actually my main question

mathslover (mathslover):

oh I helped u a lot .. hence ok

OpenStudy (lgbasallote):

and as usual...people tend to go into unnecessary details

mathslover (mathslover):

@vishweshshrimali5 is excellent in this topic but he is not online .. :(

OpenStudy (lgbasallote):

so anyway...tell me how to solve this

OpenStudy (lgbasallote):

i get how the 4.5 was got

OpenStudy (lgbasallote):

but why 4/4.5

OpenStudy (lgbasallote):

lol

OpenStudy (unklerhaukus):

so we have 3 mol A from the stoichiometry 2 mol A -> 3 mol C so we will theoretically make 3/2 *3 mol c= 9/2 mol C

OpenStudy (unklerhaukus):

yield was 4 mol C

OpenStudy (lgbasallote):

yes i got 9/2 too @UnkleRhaukus but why 4/9/2?

OpenStudy (unklerhaukus):

\[\text{relative % yield}=\frac{\text {yield}}{\text{theoretical yield}}\times100\%\]

OpenStudy (lgbasallote):

ohh yes...how could i forget

OpenStudy (unklerhaukus):

\[=\frac{4}{9/2}\times100\%=800\%/9\approx89\%\]

OpenStudy (lgbasallote):

my answer key still says 89 is wrong though

OpenStudy (unklerhaukus):

what is suggested by the key?

OpenStudy (unklerhaukus):

does the answer key's answer make sense?